Let f (x) = 5x^2– 6 and g(x) = x – 2.
(a) Find the composite function ) ( f o g)(x) and simplify. Show work.
(b) Find ( f o g )(1).

Answer by Cromlix(4381)

Hi there,
a) f(x) = 5x^2 - 6: g(x) = x - 2
(f o g((x)= 5(x - 2)^2 - 6
(f o g)(x) = 5(x^2 - 4x + 4) - 6
(f o g)(x) = 5x^2 - 20x + 20 - 6
(f o g)(x) = 5x^2 - 20x + 14
b)(f o g)(1) = 5(1)^2 - 20(1) + 14
(f o g)(1) = 5 - 20 + 14
(f o g)(1) = -1
Hope this helps :-)